absolute value of -175
absolute value of -4+10
-6-absolute value of -48

Answer by ReadingBoosters(3246)

|-175| = 175
...
|-4+10| = |6| = 6
...
-6 - |-48| = -6 - 48 = -54
.....................
Have more questions?
Email: HomeworkHelpers@readingboosters.com
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.